# Determination of the energy scale of cosmic ray measurements using the Auger Engineering Radio Array

### v1

@@ 3 Event selection and simulation production
For each of these 844 air showers we prepare one CoREAS simulation each with a proton and an iron primary according to the logic described above. The chosen interaction models are Sybill 2.3d and UrQMD. We note, however, that our result does not depend on the choice of hadronic interaction model as the radio emission is purely sensitive to the electromagnetic cascade of the air shower and we subtract the invisible energy of the chosen hadronic interacion model in a self-consistent way. For the simulations, we adopt a per-event atmospheric density and refractivity gradient [15] derived from the Global Data Assimilation System (GDAS). Furthermore, very conservative settings on particle thinning and a fine-grained treatment of multiple scattering (STEPFC parameter of 0.05) [16] ensure absolute signal predictions with the highest-possible accuracy. The simulated radio signals are then fed through a detailed detector simulation with the Offline analysis framework [13], including the addition of measured background noise taken at a given antenna station within 10 minutes of the measured event, before being reconstructed in the exact same way as the data.
@@ 5 Systematic uncertainties
Of the uncertainties related to predicting the absolute strength of the radio emission, the dominant one with 5.1% is the “radio emission yield”, i.e., the relation between the absolute radio-emission strength and the energy in the underlying electromagnetic cascade. This value results from in-depth comparisons of independent implementations of both the simulations of extensive air showers and their associated radio emission with first-principle classical electrodynamics within CORSIKA 7, ZHAireS, CORSIKA 8 [16, 17] and with the “endpoints” and ZHS formalisms [17]. In terms of experimental verification, the only measurement available so far is that of the SLAC T-510 experiment, which showed agreement between measured and simulated signals on the level of 5% with a systematic uncertainty of $\sim 10$% [18]. Again, other systematic uncertainties related to the signal prediction are negligible.
@@ 6 Conclusions
We have used a high-quality data set of 844 air showers measured with AERA in the energy range from $3\cdot 10^{17}$ eV to several $10^{18}$ eV in comparison with simulations of the same air showers using the CoREAS code to independently determine the absolute energy scale of cosmic-ray measurements at the Pierre Auger Observatory using radio signals. Particular care has been taken in crafting an analysis that does not suffer from biases and in which all individual systematic uncertainties are under control at the 5% level. The analysis was developed using a blinded data set to avoid subconscious influence on the analysis and the choice of selection cuts. The end result is that the energy scale determined from AERA is 12% higher than the one from the FD, a constant factor over the investigated energy range and the period from 2014 to 2020. The systematic uncertainty on the energy scale determined from radio data amounts to 10.3% when building on the “radio emission yield” uncertainty estimated from simulations, or 13.4% when relying on the measurement uncertainty reported by the SLAC T-510 experiment. The energy scale derived from AERA is well in agreement with the Auger energy scale determined from measurements with the Fluorescence Detector, a strong and independent confirmation of the FD energy scale within systematic uncertainties. In the future, we will use the Auger Radio Detector covering the full 3,000 km2 of the observatory to also independently determine the energy scale at the highest energies from radio measurements.

### v2

@@ 3 Event selection and simulation production
For each of these 844 air showers we prepare one CoREAS simulation each with a proton and an iron primary according to the logic described above. The chosen interaction models are Sybill 2.3d and UrQMD. We note, however, that our result does not depend on the choice of hadronic interaction model as the radio emission is purely sensitive to the electromagnetic cascade of the air shower and we subtract the invisible energy of the chosen hadronic interacion model in a self-consistent way. For the simulations, we adopt a per-event atmospheric density and refractivity gradient [15] derived from the Global Data Assimilation System (GDAS). Furthermore, particle thinning at the $10^{-6}$ level with optimized weight limitation and a fine-grained treatment of multiple scattering (STEPFC parameter of 0.05) [16] ensure absolute signal predictions with the highest-possible accuracy. The simulated radio signals are then fed through a detailed detector simulation with the Offline analysis framework [13], including the addition of measured background noise taken at a given antenna station within 10 minutes of the measured event, before being reconstructed in the exact same way as the data.
@@ 5 Systematic uncertainties
Of the uncertainties related to predicting the absolute strength of the radio emission, the dominant one with 5.0% is the “radio emission yield”, i.e., the relation between the absolute radio-emission strength and the energy in the underlying electromagnetic cascade. This value results from in-depth comparisons of independent implementations of both the simulations of extensive air showers and their associated radio emission with first-principle classical electrodynamics within CORSIKA 7, ZHAireS, CORSIKA 8 [16, 17] and with the “endpoints” and ZHS formalisms [17]. In terms of experimental verification, the only measurement available so far is that of the SLAC T-510 experiment, which showed agreement between measured and simulated signals on the level of 5% with a systematic uncertainty of $\sim 10$% [18]. Again, other systematic uncertainties related to the signal prediction are negligible.
@@ 6 Conclusions
We have used a high-quality data set of 844 air showers measured with AERA in the energy range from $3\cdot 10^{17}$ eV to several $10^{18}$ eV in comparison with simulations of the same air showers using the CoREAS code to independently determine the absolute energy scale of cosmic-ray measurements at the Pierre Auger Observatory using radio signals. Particular care has been taken in crafting an analysis that does not suffer from biases and in which all individual systematic uncertainties are under control at the 5% level. The analysis was developed using a blinded data set to avoid subconscious influence on the analysis and the choice of selection cuts. The end result is that the energy scale determined from AERA is 12% higher than the one from the FD, a constant factor over the investigated energy range and the period from 2014 to 2020. The systematic uncertainty on the energy scale determined from radio data amounts to 10.2% when building on the “radio emission yield” uncertainty estimated from simulations, or 13.4% when relying on the measurement uncertainty reported by the SLAC T-510 experiment. The energy scale derived from AERA is well in agreement with the Auger energy scale determined from measurements with the Fluorescence Detector, a strong and independent confirmation of the FD energy scale within systematic uncertainties. In the future, we will use the Auger Radio Detector covering the full 3,000 km2 of the observatory to also independently determine the energy scale at the highest energies from radio measurements.